# Alleviating Non-identifiability: a High-fidelity Calibration Objective for Financial Market Simulation with Multivariate Time Series Data

### v3

@@ Abstract
The non-identifiability issue has been frequently reported in the social simulation works, where different parameters of an agent-based simulation model yield indistinguishable simulated time series data under certain discrepancy metrics. This issue largely undermines the simulation fidelity yet lacks dedicated investigations. This paper theoretically analyzes that incorporating multiple time series data features in the model calibration phase can alleviate the non-identifiability exponentially with the increasing number of features. To implement this theoretical finding, a maximization-based aggregation function is proposed based on existing discrepancy metrics to form a new calibration objective function. For verification, the task of calibrating the Financial Market Simulation (FMS), a typical yet complex social simulation, is considered. Empirical studies confirm the significant improvements in alleviating the non-identifiability of calibration tasks. Furthermore, as a model-agnostic method, it achieves much higher simulation fidelity of the chosen FMS model on both synthetic and real market data. Hence, this work is expected to provide not only a rigorous understanding of non-identifiability in social simulation, but an off-the-shelf high-fidelity calibration objective function for FMS.
1 Introduction
@@
Social simulation plays a pivotal role of understanding the complex social systems in a way of revealing their endogenous complexity  [27]. By properly modeling the endogenous components of the social system as multiple interacting agents, i.e., using the Agent-based Models (ABMs)  [10], various “what-if” tests can be conducted by intervening the agents of interest and analyzing the evolution process of the simulators  [8, 29].
@@
The success of a social simulation largely lies with the model fidelity. Note that, the what-if tests are often counterfactual, i.e., they do not happen in real social systems and thus there are no ground-truth for validation. To what extent should we believe the results of what-if tests? Besides, due to the uncertainty of human activities, it is often impractical to derive mathematically trustworthy social simulation models. Usually, this issue is addressed in a data-driven manner via model calibration  [16].
@@
Specifically, a simulation model $M(\omega)$ is essentially a data generative process, i.e., $M(\omega)=\mathbf{X}_{T^{\prime}},T^{\prime}=[1,...,\tau],\tau\in\mathbb{N}^{+}$. If the simulated data $\mathbf{X}_{T}$ resembles the observed data $\hat{\mathbf{X}}_{T}$ of the targeted real system within any discrete time interval of interest $T=[t_{s},t_{s}+1,...,t_{e}]$, it is considered that the underlying data generating probabilistic distribution of the real system for generating $\hat{\mathbf{X}}_{T}=\{\hat{\mathbf{X}}_{t}\}^{t_{e}}_{t=t_{s}}$ is learned by $M(\omega)$. Hence, while using $M(\omega)$ for what-if test, by keeping the non-intervened agents of $M(\omega)$ unchanged, the simulation results of how the intervened agent influences the system are considered trustworthy. To obtain such $M(\omega)$, the discrepancy between $\mathbf{X}_{T}$ and $\hat{\mathbf{X}}_{T}$ is measured by various metrics $D$  [24]. The smaller $D$ is, the closer $M(\omega)$ approximates the data generating distribution of the real social system. By fixing the ABM-based model structure $M$, calibrating the parameter $\omega$ to fit a given observed data sequence $\hat{\mathbf{X}}_{T}$ is often modeled as an optimization problem  [24]:
$\min_{\omega\in\Omega^{D}}D(\hat{\mathbf{X}}_{T},M(\omega)=\mathbf{X}_{T})$ (1)
@@
In practice, the calibration methods often suffer from the so-called non-identifiability issue  [11, 1]. That is, there are multiple parameters in the parameter space $\Omega^{D}$ share the same discrepancy value to the given $\hat{\mathbf{X}}_{T}$. However, those parameters highly likely lead to different distributions of data generation, due to the nonlinear model structure $M(\omega)$. In other words, the “ground-truth” parameter of deciding the “ground-truth” data generating distribution cannot be effectively identified from the other parameters in those sets. Consequently, the what-if tests on the resultant model of a randomly picked non-identifiable parameter therefrom are less trustworthy.
@@
Despite that the non-identifiability is frequently mentioned in the literature, there is neither a formal definition to describe it nor a rigorous way to alleviate it. This paper first defines the non-identifiability as the probability of a randomly uniformly sampled parameter falling into the non-identifiable set. With this definition, this paper mathematically proves that the non-identifiability can be alleviated by calibrating with more distinct features, each of which represents an observed univariate time series. More specifically, we mathematically show that the upper bound of the non-identifiability can be reduced exponentially with the number of features, if the selected features satisfy certain conditions.
@@
This explains well that the previous works mostly calibrate their models to merely one data feature and is thus less sufficient to describe the uniqueness of a social system. For clarification and applicability, this work restricts the scenarios to the Financial Market Simulation (FMS), as it is a representative social simulation and has been active for over 30 years  [22]. Like other complex social systems, the real-world financial markets will output multivariate states data over time, e.g., price, volume, bid/ask directions, order arrival time, and other hand-crafted features  [21]. However, existing works merely consider the observed mid-price data for calibration, resulting in wide criticism for their low simulation fidelity  [9], especially in the high-frequency intraday simulations  [25, 11].
@@
Note that, following our theory, the task of calibrating to multivariate time series is naturally divided into multiple individual tasks, each of which calibrates the model to one distinct univariate time series. In practice, how to jointly execute those individual tasks so that the obtained parameter is more identifiable to the original calibration tasks with multivariate time series? This challenges the implementation of the above theory. To address it, we propose a novel calibration objective by aggregating all those individual calibration tasks using the maximization function. This new objective function mathematically asks to search in the intersection of all those individual spaces, thus the obtained parameter can be highly identifiable across all individual tasks and thus fit our theory accurately.
@@
Extensive empirical studies have successfully shown that the proposed method can significantly alleviate the non-identifiability issue of financial market simulation. Such advantages are verified robust over 6 commonly seen financial market data features. Besides, this paper also presents the high-fidelity simulation results by using the proposed new calibration objective function on 10 synthetic data and 1 real data from Shenzhen Stock Exchange of China.
@@
The reminder of this paper is as follows. Section 2 describes the background of financial market simulation and related works. Section 3 theoretically discuss the alleviation of non-identifiability. Section 4 reports the empirical studies in details. Section 5 concludes this work.
2 Background
2.1 Calibration of Social Simulators
@@
To ensure the simulation fidelity, the simulator often needs careful calibration of the agent parameters, so that the simulated data resembles the observed time series from the real system. The general form of calibration refers to Eq. (1).
@@
In recent years, the study of the discrepancy metrics $D$ has attracted increasing research attentions, as the distance measure between time series remains unsolved and impacts significantly on the calibration performance. Methods of Simulated Moments  [12] are a typical method that first transforms the time series into several statistical moments, and then calculates the weighted average between the moments of observed data and simulated data. The information criterion based objective functions  [18, 2] abstract the temporal information of the original time series via various techniques like histograms and measure the distances between the abstracted temporal vectors. The non-parametric Kolmogorov-Smirnov test is also introduced as $D$ by statistically testing between the probabilistic distributions of the observed data and the simulated data  [1]. Other methods employ the Bayesian theory to estimate the likelihood  [17] or posterior distribution  [13] of the observed data also require a well-defined $D$ for selecting effective samples and updating the estimated distributions.
@@
Despite of the progresses of exploiting various metrics for calibrating social simulators, existing works frequently encounter the so-called non-identifiability issue, resulting in poor simulation fidelity  [19]. From our viewpoint, one major reason is that existing works merely use the one single feature to calibrate the model, ignoring that the real systems continuously output multivariate data. Although the used feature might be a major factor to describe the social system, the other features can help capture more details of the system. Unfortunately, existing calibration objectives are mostly designed for univariate time series data. Consequently, existing works are ill-equipped to calibrate with multivariate time series data. Yet, it is theoretically unclear how exactly it benefits from calibrating more features.
2.2 ABM-based FMS and Data Features
In social simulations, ABMs are dominantly preferred than the black-box deep learning models due to their interpretability of the data generating process  [14, 20]. ABM constructs a collection of agents mimicking various social activities and interacting together to simulate the whole social system  [23]. In FMS, the agents are designed to model various trading strategies  [4] (illustrated in Figure 1), e.g., fundamentalists, momentum traders, high-frequency traders, market makers, and arbitrage traders.
Regardless of the differences among agents, the ABM-based FMSs follow the same workflow. That is, at the beginning of the simulation, the parameterized agents are initialized with the calibrated parameters, which determine how exactly those agents submit orders under their own simulated strategies. Each submitted order usually consists of the price, volume, direction, and arrival time. The Limited Order Book (LOB), a special data structure for organizing the existing untraded orders, is maintained by the exchange and can be observed by all trading agents. The simulated exchange generates the next LOB by matchmaking the incoming orders submitted by the agents and the current LOB  [28].
Those untraded orders are categorized into two directions (bid or ask) and sorted by their prices. For the orders with the same price, they are again sorted by the arrival time. Each agent continuously observes the current LOB and submit its order at will. If a new order matches a price in the opposite direction of the LOB, it will be traded immediately, and the matched orders will be removed from the LOB. Undoubtedly, many data features can be extracted from LOB  [21]. However, existing FMSs mostly considered the mid-price as the only feature for calibration, which represents the average between the best ask price and best bid prices (see Figure 2) and implies the potential price movement of the market.
@@ 2.2 ABM-based FMS and Data Features
3 A High-fidelity Calibration Objective
@@
3.1 The Definition of Non-identifiability
As existing discrepancy metrics $D$ are designed for univariate data, we re-denote the target data as $\hat{\mathbf{X}}_{T}^{k}$ with a superscript $k$ to emphasize that it is a 1-dimensional time series describing some $k$-th feature of the real system. In financial markets, this feature can be either price, volume, bid/ask directions, order arrival time, or any other hand-crafted features  [21]. Correspondingly, the $k$-th feature of the simulated data is re-denoted as $\mathbf{X}^{k}_{T}$. Let $\Omega^{D,k}$ be the parameter space of this calibration task. Then the original calibration task can be re-written as
$\min_{\omega\in\Omega^{D,k}}D(\hat{\mathbf{X}}_{T}^{k},M(\omega)=\mathbf{X}^{k}_{T}).$ (2)
Without loss of generality, it is assumed that the optimal parameter $\omega^{*}$ lies in the defined parameter space, i.e., $\omega^{*}\in\Omega^{D,k}$, where $D(\hat{\mathbf{X}}_{T}^{k},M(\omega^{*}))<\epsilon$. Note that, the optimal parameter may not be unique by definition. The cardinality of the optimal set is influenced by $\epsilon$, which is used to filter the observation noise of data. Then the non-identifiable set is defined as follow:
@@ 3.1 The Definition of Non-identifiability
There exists some region $\mathcal{S}^{D,k,\epsilon}\subseteq\Omega^{D,k}$ that for all $\omega_{1},\omega_{2}\in\mathcal{S}^{D,k,\epsilon}$ such that $D(\hat{\mathbf{X}}_{T}^{k},M(\omega_{1}))<\epsilon$ and $D(\hat{\mathbf{X}}_{T}^{k},M(\omega_{2}))<\epsilon$. Such region $\mathcal{S}^{D,k,\epsilon}$ is called the non-identifiable set.
The cardinality of $\mathcal{S}^{D,k,\epsilon}$, denoted as $|\mathcal{S}^{D,k,\epsilon}|$, is also influenced by $\epsilon$. Then the non-identifiability can be defined as follows:
@@
Let $P(\omega\in\mathcal{S}^{D,k,\epsilon}|\hat{\mathbf{X}}_{T}^{k})$ be the non-identifiability to the $k$-th univariate time series data. It it defined as the probability of uniformly randomly sampling a parameter $\omega$ from $\Omega^{D,k}$ while it falls into $\mathcal{S}^{D,k,\epsilon}$. And we have
$P\left(\omega\in\mathcal{S}^{D,k,\epsilon}\middle|\hat{\mathbf{X}}_{T}^{k}\right)=\frac{|\mathcal{S}^{D,k,\epsilon}|}{|\Omega^{D,k}|}.$ (3)
Therefore, the rank of any $\omega_{1},\omega_{2}\in\mathcal{S}^{D,k,\epsilon}$ cannot be effectively identified by the objective $D$ given the data $\hat{\mathbf{X}}_{T}^{k}$, leading to that the comparison-based calibration process easily gets stuck in $\mathcal{S}^{D,k,\epsilon}$. Note that, $\mathcal{S}^{D,k,\epsilon}$ can involve arbitrary parameters in $\Omega^{D,k}$, including the “ground-truth” parameter $\omega^{*}$. Hence, if the calibration process converges to $\mathcal{S}^{D,k,\epsilon}$, one has to randomly pick one $\omega\in\mathcal{S}^{D,k,\epsilon}$ for further what-if simulation. Unfortunately, the underlying data generative distribution decided by any randomly picked $\omega\in\mathcal{S}^{D,k,\epsilon}$ is less likely to be the same with $\omega^{*}$, given the nonlinear nature of the simulation model $M(\omega)$. Consequently, the what-if test on the randomly picked $\omega$ is less trustworthy, unless the cardinality of $\mathcal{S}^{D,k,\epsilon}$ is sufficiently small.
3.2 Alleviating the Non-identifiability with Multiple Features

### v7

@@ Abstract
The non-identifiability issue has been frequently reported in social simulation works, where different parameters of an agent-based simulation model yield indistinguishable simulated time series data under certain discrepancy metrics. This issue largely undermines the simulation fidelity yet lacks dedicated investigations. This paper theoretically demonstrates that incorporating multiple time series data features during the model calibration phase can exponentially alleviate non-identifiability as the number of features increases. To implement this theoretical finding, a maximization-based aggregation function is proposed based on existing discrepancy metrics to form a new calibration objective function. For verification, the task of calibrating the Financial Market Simulation (FMS), a typical yet complex social simulation, is considered. Empirical studies confirm the significant improvements in alleviating the non-identifiability of calibration tasks. Furthermore, as a model-agnostic method, it achieves much higher simulation fidelity of the chosen FMS model on both synthetic and real market data. Moreover, it is both theoretically and empirically analyzed that as long as the features are selected and not linearly correlated, they can contribute to alleviation, which demonstrates the robustness of the proposed objective. Hence, this work is expected to provide not only a rigorous understanding of non-identifiability in social simulation but also an off-the-shelf high-fidelity calibration objective function for FMS.
Index Terms:
Financial market simulation, Agent-based modeling, Non-identifiability, High-fidelity calibration objective
I Introduction
@@
Social simulation plays a pivotal role in understanding complex social systems in a way of revealing their endogenous complexity [1]. By properly modeling the endogenous components of the social system as multiple interacting agents, i.e., using the Agent-based Models (ABMs) [2], various what-if tests can be conducted by intervening the agents of interest and analyzing the evolution process of the simulators [3, 4]. These what-if tests involve simulating hypothetical scenarios to assess how changes in agent behaviors or interactions may influence the overall system dynamics.
@@
The success of a social simulation largely lies in the model’s fidelity. Note that the what-if tests are often counterfactual, i.e., they do not happen in real social systems, and thus, there are no grounds for validation. To what extent should we believe the results of what-if tests? Besides, due to the uncertainty of human activities, it is often impractical to derive mathematically trustworthy social simulation models. Usually, this issue is addressed in a data-driven manner via model calibration [5].
@@
Specifically, a simulation model $M(\omega)$ is essentially a data-generating process that can produce sequences of arbitrary length, i.e., $M(\omega)=\mathbf{X}_{T^{\prime}}$, where $T^{\prime}=[1,...,\tau]$ and $\tau\in\mathbb{N}^{+}$. If the simulated data $\mathbf{X}_{T}$ resembles the observed data $\hat{\mathbf{X}}_{T}$ of the targeted real system within any discrete time interval of interest $T=[t_{s},t_{s}+1,...,t_{e}]$, we consider that $M(\omega)$ has learned the underlying data-generating probabilistic distribution of the real system for generating $\hat{\mathbf{X}}_{T}$. Therefore, when using $M(\omega)$ for what-if tests, by keeping the non-intervened agents of $M(\omega)$ unchanged, the simulation results of how the intervened agent influences the system are considered trustworthy. To obtain such $M(\omega)$, the discrepancy between $\mathbf{X}_{T}$ and $\hat{\mathbf{X}}_{T}$ is measured by various metrics $D$ [6]. The smaller $D$ is, the closer $M(\omega)$ approximates the data-generating distribution of the real social system. By fixing the ABM-based model structure $M$, calibrating the parameter $\omega$ to fit a given observed data sequence $\hat{\mathbf{X}}_{T}$ is often modeled as an optimization problem [6, 7]:
$\min_{\omega\in\Omega^{D}}D\left(\hat{\mathbf{X}}_{T},M\left(\omega\right)=\mathbf{X}_{T}\right).$ (1)
@@
In practice, the calibration methods often suffer from the so-called non-identifiability issue [8, 9]. That is, there are multiple parameters in the parameter space $\Omega^{D}$ share the same discrepancy value to the given $\hat{\mathbf{X}}_{T}$. However, those parameters highly likely lead to different distributions of data generation due to the nonlinear model structure $M(\omega)$. In other words, the “ground-truth” parameter of deciding the “ground-truth” data generating distribution cannot be effectively identified from the other parameters in those sets. Consequently, the what-if tests on the resultant model with a randomly picked non-identifiable parameter from there are less trustworthy.
Take the simulation of the well-known flash crush as an example. Two cases of non-identifiability are illustrated in Fig. 1. In case 1, the two calibrated parameters represent different trading behaviors while both two corresponding simulated data enjoy identical objective values. By analyzing the flash crush with the two parameters, one may get different causes (i.e., high liquidity taking with a large volume of market orders v.s. potential manipulation with a large volume of canceling orders). Also, the non-identifiability can lead to case 2, where two simulated data have the same discrepancy to the target data but appear to be totally different trends. In both cases, the calibration algorithm cannot identify the one with higher fidelity, which can deteriorate the downstream analysis.
Figure 1: Two cases of non-identifiable simulation.
@@
Even though non-identifiability is frequently mentioned in the literature, there is neither a formal definition to describe it nor a rigorous way to alleviate it. This paper first defines non-identifiability as the probability of a randomly uniformly sampled parameter falling into the non-identifiable set. With this definition, this paper mathematically proves that the non-identifiability can be alleviated by calibrating with more distinct features, each of which represents an observed univariate time series. More specifically, we mathematically show that the upper bound of the non-identifiability can be reduced exponentially with the number of features if the selected features satisfy certain conditions.
@@
This explains well that the previous works mostly calibrate their models to merely one data feature and are thus less sufficient to describe the uniqueness of a social system. For clarification and applicability, this work restricts the scenarios to the FMS, as it is a representative social simulation and has been active for over 30 years [10]. Like other complex social systems, the real-world financial markets will output multivariate state data over time, e.g., price, volume, bid/ask directions, order arrival time, and other handcrafted features [11]. However, existing works merely consider the observed mid-price data for calibration, resulting in wide criticism for their low simulation fidelity [12], especially in the high-frequency intraday simulations [13, 8].
@@
Note that, following our theory, the task of calibrating to a multivariate time series is naturally divided into multiple individual tasks, each of which calibrates the model to one distinct univariate time series. In practice, how to jointly execute those individual tasks so that the obtained parameter is more identifiable to the original calibration tasks with multivariate time series? This challenges the implementation of the above theory. To address it, we propose a novel calibration objective by aggregating all those individual calibration tasks using the maximization function. This new objective function mathematically asks to search in the intersection of all those individual spaces. Thus, the obtained parameter can be highly identifiable across all individual tasks and thus fit our theory accurately.
@@
Extensive empirical studies have successfully shown that the proposed method can significantly alleviate the non-identifiability issue of FMS. Such advantages are verified robust over 6 commonly seen features of financial market data. Besides, this paper also presents the high-fidelity simulation results by using the proposed new calibration objective function on 10 synthetic data and 1 real data from the Shenzhen Stock Exchange of China.
@@
The remainder of this paper is as follows. Section II describes the background of FMS and related works. Section III theoretically discusses the alleviation of non-identifiability. Section IV reports the empirical studies in detail. Section V concludes this work.
II Background
II-A Calibration of Social Simulators
@@
To ensure the simulation fidelity, the simulator often needs careful calibration of the agent parameters so that the simulated data resembles the observed time series from the real system. The general form of calibration refers to (1).
@@
In recent years, the study of the discrepancy metrics $D$ has attracted increasing research attention, as the distance measure between time series remains unsolved and impacts significantly on the calibration performance. Methods of Simulated Moments [14] is a typical method that first transforms the time series into several statistical moments and then calculates the weighted average between the moments of observed data and simulated data. The information criterion-based objective functions [15, 16] abstract the temporal information of the original time series via various techniques like histograms and measure the distances between the abstracted temporal vectors. The non-parametric Kolmogorov-Smirnov test is also introduced as $D$ by statistically testing between the probabilistic distributions of the observed data and the simulated data [9]. Other methods employ the Bayesian theory to estimate the likelihood [17] or posterior distribution [18] of the observed data also require a well-defined $D$ for selecting effective samples and updating the estimated distributions.
@@
Despite the progress of exploiting various metrics for calibrating social simulators, existing works frequently encounter the so-called non-identifiability issue, resulting in poor simulation fidelity [19]. From our viewpoint, one major reason is that existing works merely use one single feature to calibrate the model, ignoring that the real systems continuously output multivariate data. Although the features used might be a major factor in describing the social system, other features can help capture more details of the system. Unfortunately, existing calibration objectives are mostly designed for univariate time series data. Consequently, existing works are ill-equipped to calibrate with multivariate time series data. Yet, it is theoretically unclear how exactly it benefits from calibrating more features.
II-B ABM-based FMS and Data Features
In social simulations, ABMs are dominantly preferred over the black-box deep learning models due to their interpretability of the data-generating process [20, 21]. ABM constructs a collection of agents mimicking various social activities and interacting together to simulate the whole social system [22]. In FMS, the agents are designed to model various trading strategies [23] (illustrated in Fig. 2), e.g., fundamentalists, momentum traders, high-frequency traders, market makers, and arbitrage traders.
@@ II-B ABM-based FMS and Data Features
Regardless of the differences among agents, the ABM-based FMSs follow the same workflow [24]. That is, at the beginning of the simulation, the parameterized agents are initialized with the calibrated parameters, which determine how exactly those agents submit orders under their own simulated strategies. Each submitted order usually consists of the price, volume, direction, and arrival time. The Limited Order Book (LOB), a particular data structure for organizing the existing untraded orders, is maintained by the exchange and can be observed by all trading agents. The simulated exchange generates the next LOB by matchmaking the incoming orders submitted by the agents and the current LOB [25].
Those untraded orders are categorized into two directions (bid or ask) and sorted by their prices. Orders with the same price are again sorted by arrival time. Each agent continuously observes the current LOB and submits its order at will. If a new order matches a price in the opposite direction of the LOB, it will be traded immediately, and the matched orders will be removed from the LOB. Undoubtedly, many data features can be extracted from LOB [11]. However, existing FMSs mostly considered the mid-price as the only feature for calibration, which represents the average between the best ask price and best bid prices (see Fig. 3) and implies the potential price movement of the market.
Recent advances in financial time series modeling prioritize two objectives: capturing nonlinear dynamics while maintaining interpretability. The RHINE model [26] addresses regime-switching through kernel-based spectral clustering and sliding-window eigengap analysis. Parallel innovations in concept drift detection, exemplified by WormKAN [27], employ Kolmogorov-Arnold Networks (KANs) with self-representation matrices to track abrupt transitions. These are unified through MT-KAN’s [28] lightweight architectures for multivariate interaction modeling. Besides that, recent works like [9] and [21] tried to calibrate simulators with multi-variate time series but did not investigate the non-identifiability issue from the perspective of using more features.
III A High-fidelity Calibration Objective
@@
III-A The Definition of Non-identifiability
As existing discrepancy metrics $D$ are designed for univariate data, we re-denote the target data as $\hat{\mathbf{X}}_{T}^{k}$ with a superscript $k$ to emphasize that it is a 1-dimensional time series describing some $k$-th feature of the real system. In financial markets, this feature can be either price, volume, bid/ask directions, order arrival time, or any other handcrafted features [11]. Correspondingly, the $k$-th feature of the simulated data is re-denoted as $\mathbf{X}^{k}_{T}$. Let $\Omega^{D,k}$ be the parameter space of this calibration task. Then the original calibration task can be re-written as
$\min_{\omega\in\Omega^{D,k}}D\left(\hat{\mathbf{X}}_{T}^{k},M(\omega)=\mathbf{X}^{k}_{T}\right).$ (2)
Without loss of generality, it is assumed that the optimal parameter $\omega^{*}$ lies in the defined parameter space, i.e., $\omega^{*}\in\Omega^{D,k}$, where $D(\hat{\mathbf{X}}_{T}^{k},M(\omega^{*}))<\epsilon$. Note that the optimal parameter may not be unique by definition. The cardinality of the set containing all optimal parameters depends on $\epsilon$, which filters the observation noise in the data. This set, referred to as the non-identifiable set, is formally defined as follows:
@@ III-A The Definition of Non-identifiability
There exists some region $\mathcal{S}^{D,k,\epsilon}\subseteq\Omega^{D,k}$ that for all $\omega_{1},\omega_{2}\in\mathcal{S}^{D,k,\epsilon}$ $D(\hat{\mathbf{X}}_{T}^{k},M(\omega_{1}))<\epsilon$ and $D(\hat{\mathbf{X}}_{T}^{k},M(\omega_{2}))<\epsilon$. Such region $\mathcal{S}^{D,k,\epsilon}$ is called the non-identifiable set.
The Lebesgue measure (intuitively the hypervolume) of $\mathcal{S}^{D,k,\epsilon}$, denoted as $\mu\left(\mathcal{S}^{D,k,\epsilon}\right)$, is influenced by $\epsilon$. By choosing the parameter space to be finite and based on Definition 1, we know that $0<\mu\left(\mathcal{S}^{D,k,\epsilon}\right)\leq\mu\left(\Omega^{D,k,\epsilon}\right)<+\infty$. Then the non-identifiability can be defined as follows:
@@
Let $P\left(\omega\in\mathcal{S}^{D,k,\epsilon}\middle|\hat{\mathbf{X}}_{T}^{k}\right)$ be the non-identifiability to the $k$-th univariate time series data. It is defined as the probability of uniformly randomly sampling a parameter $\omega$ from $\Omega^{D,k}$ while it falls into $\mathcal{S}^{D,k,\epsilon}$. And we have
$P\left(\omega\in\mathcal{S}^{D,k,\epsilon}\middle|\hat{\mathbf{X}}_{T}^{k}\right)=\frac{\mu\left(\mathcal{S}^{D,k,\epsilon}\right)}{\mu\left(\Omega^{D,k}\right)}.$ (3)
Therefore, the rank of any $\omega_{1},\omega_{2}\in\mathcal{S}^{D,k,\epsilon}$ cannot be effectively identified by the objective $D$ given the data $\hat{\mathbf{X}}_{T}^{k}$, leading to that the comparison-based calibration process easily gets stuck in $\mathcal{S}^{D,k,\epsilon}$. Note that, $\mathcal{S}^{D,k,\epsilon}$ can involve arbitrary parameters in $\Omega^{D,k}$, including the “ground-truth” parameter $\omega^{*}$. Hence, if the calibration process converges to $\mathcal{S}^{D,k,\epsilon}$, one has to randomly pick one $\omega\in\mathcal{S}^{D,k,\epsilon}$ for further what-if simulation. Unfortunately, the underlying data-generating distribution decided by any randomly picked $\omega\in\mathcal{S}^{D,k,\epsilon}$ is less likely to be the same with $\omega^{*}$, given the nonlinear nature of the simulation model $M(\omega)$. Consequently, the what-if test on the randomly picked $\omega$ is less trustworthy unless the cardinality of $\mathcal{S}^{D,k,\epsilon}$ is sufficiently small.
III-B Alleviating the Non-identifiability with Multiple Features